# On obstructions to the Euler system method for Rankin-Selberg convolutions

### v2

@@ 4 Examples and counter-examples
The quotient map $\rho_{g}(G_{{\mathbb{Q}}})/\rho_{g}(G_{H})\rightarrow\tilde{\rho_{g}}(G_{{\mathbb{Q}}})/\tilde{\rho}_{g}(G_{H})$ is surjective between groups of same cardinality, so it is injective. Therefore, $\rho_{g}(G_{{\mathbb{Q}}})$ and $\rho_{g}(G_{H})$ have the same scalar matrices. Any pre-image by $\rho_{g}$ of any scalar matrix lies in $G_{K}$, thus $\rho_{g}(G_{{\mathbb{Q}}})$ and $\rho_{g}(G_{HK})$ contain the same scalar matrices. Because $i\notin\varepsilon_{g}(G_{HK})$, $\rho_{g}(G_{HK})$ does not contain a scalar matrix with order $8$, so neither does $\rho_{g}(G_{{\mathbb{Q}}})$. Thus, the scalar matrices in $\rho_{g}(G_{{\mathbb{Q}}})$ are exactly ±$\pm I_{2},\pm iI_{2}$.
Next, we show that there are elements $N_{-},N_{4}\in\rho_{g}(G_{H_{0}})\backslash\rho_{g}(G_{H})$ whose image in $\operatorname{PGL}_{2}(L_{0})$ has order $2$ such that $\det{N_{-}}=-1$ and $\det{N_{4}}$ has order $4$. For $N_{-}$, we can simply take the complex conjugation.
@@
The determinant map $\delta:\tilde{\rho}_{g}(G_{{\mathbb{Q}}})\rightarrow\varepsilon_{g}(G_{{\mathbb{Q}}})/\{\pm 1\}$ is onto a group of cardinality $2$. Then $\delta$ vanishes on the complex conjugation, so $\delta$ does not vanish on the cyclic subgroup $\tilde{\rho}_{g}(G_{K})$ of $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$. This cyclic subgroup has order $2n$ ($n$ being odd), so it contains a unique element $\kappa$ of order $2$, and $\delta(\kappa)$ is non-trivial. Because $K\not\subset H$ and $\tilde{\rho}_{g}(G_{H})\neq\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$, $\tilde{\rho}_{g}(G_{K})\cap\tilde{\rho}_{g}(G_{H})$ is the subgroup of index $2$ of $\tilde{\rho}_{g}(G_{K})$, so it has odd cardinality and does not contain $\kappa$. As such, if $N_{4}\in\rho_{g}(G_{{\mathbb{Q}}})$ is any element lifting $\kappa$, then $N_{4}\notin\rho_{g}(G_{H})$, the projective image of $N_{4}$ has order $2$, and $\det{N_{4}}$ has order $4$.
Fix an element $\tau\in G_{H_{0}}$ with $\rho_{g}(\tau)\notin\rho_{g}(G_{H})$. By the second part of Proposition 2, we can choose such an element $\tau$ in $G_{H_{0}(\mu_{p^{\infty}})}$. Let $\alpha\in L^{\times}\cap\mathcal{O}_{L,\mathfrak{p}}^{\times}$ be the element attached to $\tau$ by Corollary 2. Then $\alpha^{2}\in\mathcal{O}_{F,\mathfrak{p}}^{\times}$ and $\alpha\notin F$.
@@
Let $\mathfrak{p}\subset L_{0}$ be a good prime for $(f,g)$ of residue characteristic $p$. We will show that the representation $T_{f,g,\mathfrak{p}}$ satisfies (sE) by distinguishing between the following three cases:
@@
In case (i), pick $\sigma_{1}\in G_{H}\backslash G_{K}$. Since $g$ has CM by $K$, $\rho_{g}(\sigma_{1})$ has trace zero and determinant ±$\pm i$, so its eigenvalues are $u,-u$ for some primitive eighth root of unity $u\in L_{0}$. We can pick a square root $t$ of $2$ in $L_{0}$ such that $tu-1$ is a fourth root of unity, then $t\in(L_{0})_{\mathfrak{p}}$ is an element of its subfield $F_{\mathfrak{p}}$. By Proposition 2, we can choose $\sigma\in G_{H(\mu_{p^{\infty}})}$ such that $\rho_{g}(\sigma)=\rho_{g}(\sigma_{1})$ and the characteristic polynomial of $\sigma$ acting on $T_{f,\mathfrak{p}}$ is given by $x^{2}-tx+1$. Then the eigenvalues of $\sigma$ acting on $T_{f,g,\mathfrak{p}}$ (with ring of coefficients $\mathcal{O}_{(L_{0})_{\mathfrak{p}}}$) are given by the quadruple $(u\cdot u,u^{-1}u,u(-u),u^{-1}(-u))=(u^{2},1,-u^{2},-1)$, so $1$ is a simple eigenvalue for the actions of the action of $\sigma$ on $T_{f,g,\mathfrak{p}}$ and $T_{f,g,\mathfrak{p}}/\mathfrak{p}$, so it fulfills the condition (sE).
@@
In case (iii), let $u\in L_{0}$ be a primitive eighth root of unity such that $u,-u$ are the two eigenvalues of $N_{4}$. Then $\alpha u+\alpha u^{-1}=\pm\sqrt{2}\alpha\in(L_{0})_{\mathfrak{p}},(\alpha u)(\alpha u^{-1})=\alpha^{2}\in L_{0}$ are elements of $\mathcal{O}_{F_{\mathfrak{p}}}$, so there is a matrix $M\in\operatorname{GL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}})$ with determinant $\alpha^{2}$ such that its eigenvalues are $\alpha u,\alpha u^{-1}$. By Corollary 2, there exists $\sigma\in\tau G_{H(\mu_{p^{\infty}})}$ such that $\sigma$ acts on $T_{f,\mathfrak{p}}$ by the matrix $\alpha^{-1}M$ (in a suitable basis) and $\rho_{g}(\sigma)=N_{4}$. Then the eigenvalues of $\sigma$ acting on $T_{f,g,\mathfrak{p}}$ are given as in case (i) by $u^{2},1,-u^{2},-1$, all of which are distinct mod $\mathfrak{p}$, and we are done.
@@
In this case, the LMFDB says that $L={\mathbb{Q}}(\sqrt{3}),F={\mathbb{Q}},H={\mathbb{Q}}(\sqrt{-3})$, $f$ does not have complex multiplication, and $\varepsilon_{g}$ is a character of conductor $5$ and order $4$. Moreover, $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is the dihedral group of order $12=4\cdot 3$, and $g$ has complex multiplication by ${\mathbb{Q}}(\sqrt{-15})$, which is not contained in $H$. Furthermore, the number field LMFDB data shows that $\ker{\tilde{\rho}_{g}}\subset G_{H}$, so $[\tilde{\rho}_{g}(G_{{\mathbb{Q}}}):\tilde{\rho}_{g}(G_{H})]=[\rho_{g}(G_{{\mathbb{Q}}}):\rho_{g}(G_{H})]=2$. Since $a_{139}(g)=\pm 2i$, $iI_{2}\in\rho_{g}(G_{H})\subset\rho_{g}(G_{{\mathbb{Q}}})$. Since the coefficient field of $g$ is ${\mathbb{Q}}(i,\sqrt{6})$, it does not contain a primitive $8$-th root of unity, so the scalar matrices of $\rho_{g}(G_{{\mathbb{Q}}})$ are exactly ±$\pm iI_{2},\pm I_{2}$. Finally, $\varepsilon_{g}(G_{H})=\varepsilon_{g}(G_{{\mathbb{Q}}})=\{\pm i,\pm 1\}$, since $\varepsilon_{g}$ and $H$ have coprime conductors.

### v3

@@ 4 Examples and counter-examples
The quotient map $\rho_{g}(G_{{\mathbb{Q}}})/\rho_{g}(G_{H})\rightarrow\tilde{\rho}_{g}(G_{{\mathbb{Q}}})/\tilde{\rho}_{g}(G_{H})$ is surjective between groups of same cardinality, so it is injective. Therefore, $\rho_{g}(G_{{\mathbb{Q}}})$ and $\rho_{g}(G_{H})$ have the same scalar matrices. Any pre-image by $\rho_{g}$ of any scalar matrix lies in $G_{K}$, thus $\rho_{g}(G_{{\mathbb{Q}}})$ and $\rho_{g}(G_{HK})$ contain the same scalar matrices. Because $i\notin\varepsilon_{g}(G_{HK})$, $\rho_{g}(G_{HK})$ does not contain a scalar matrix with order $8$, so neither does $\rho_{g}(G_{{\mathbb{Q}}})$. Thus, the scalar matrices in $\rho_{g}(G_{{\mathbb{Q}}})$ are exactly $\pm I_{2},\pm iI_{2}$.
@@
Next, we show that there is $N_{4}\in\rho_{g}(G_{H_{0}})\backslash\rho_{g}(G_{H})=\rho_{g}(G_{{\mathbb{Q}}})\backslash\rho_{g}(G_{H})$ whose image in $\operatorname{PGL}_{2}(L_{0})$ has order $2$ such that $\det{N_{4}}$ has order $4$. Indeed, let $\delta:\tilde{\rho}_{g}(G_{{\mathbb{Q}}})\rightarrow\varepsilon_{g}(G_{{\mathbb{Q}}})/\{\pm 1\}$ be the determinant map: it is onto a group of cardinality $2$. Then $\delta$ vanishes on the complex conjugation, so $\delta$ does not vanish on the cyclic subgroup $\tilde{\rho}_{g}(G_{K})$ of $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$. This cyclic subgroup has order $2n$ (and $n$ is odd), so it contains a unique element $\kappa$ of order $2$, and $\delta(\kappa)$ is non-trivial. Because $K\not\subset H$ and $\tilde{\rho}_{g}(G_{H})\neq\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$, $\tilde{\rho}_{g}(G_{K})\cap\tilde{\rho}_{g}(G_{H})$ is the subgroup of index $2$ of $\tilde{\rho}_{g}(G_{K})$, so it has odd cardinality and does not contain $\kappa$. Thus, if $N_{4}\in\rho_{g}(G_{{\mathbb{Q}}})$ is any element lifting $\kappa$, then $N_{4}\notin\rho_{g}(G_{H})$, the projective image of $N_{4}$ has order $2$, and $\det{N_{4}}$ has order $4$.
@@
By the second part of Proposition 2, we can fix an element $\tau\in G_{H_{0}(\mu_{p^{\infty}})}$ such that $\rho_{g}(\tau)\notin\rho_{g}(G_{H})$. Let $\alpha\in L^{\times}\cap\mathcal{O}_{L,\mathfrak{p}}^{\times}$ be the element attached to $\tau$ by Corollary 2. Then $\alpha^{2}\in\mathcal{O}_{F,\mathfrak{p}}^{\times}$ and $\alpha\notin F$. Let $\mathfrak{p}\subset\mathcal{O}_{L_{0}}$ be a good prime for $(f,g)$ of residue characteristic $p$. We will show that the representation $T_{f,g,\mathfrak{p}}$ satisfies (sE) by distinguishing between the following three cases:
@@
[…]
@@
In case (iii), let $u\in L_{0}$ be a primitive eighth root of unity such that u,$u$,-u are the two eigenvalues of $N_{4}$. Then the two elements $\alpha u+\alpha u^{-1}=\pm\sqrt{2}\alpha,(\alpha u)(\alpha u^{-1})=\alpha^{2}$ of $(L_{0})_{\mathfrak{p}}$ lie in the subring $\mathcal{O}_{F_{\mathfrak{p}}}$, so there exists a matrix $M\in\operatorname{GL}_{2}(\mathcal{O}_{F_{\mathfrak{p}}})$ with eigenvalues $\alpha u,\alpha u^{-1}$, so that $\det{M}=\alpha^{2}$. By Corollary 2, there exists $\sigma\in\tau G_{H(\mu_{p^{\infty}})}$ acting on $T_{f,\mathfrak{p}}$ by the matrix $\alpha^{-1}M$ (in a suitable basis) and such that $\rho_{g}(\sigma)=N_{4}$. Then the eigenvalues of $\sigma$ acting on $T_{f,g,\mathfrak{p}}$ are given as in case (i) by $u^{2},1,-u^{2},-1$, all of which are distinct mod $\mathfrak{p}$, and we are done.
@@
In this case, the LMFDB says that $L={\mathbb{Q}}(\sqrt{3}),F={\mathbb{Q}},H={\mathbb{Q}}(\sqrt{-3})$, $f$ does not have complex multiplication, and $\varepsilon_{g}$ is a character of conductor $5$ and order $4$. Moreover, $\tilde{\rho}_{g}(G_{{\mathbb{Q}}})$ is the dihedral group of order $12=4\cdot 3$, and $g$ has complex multiplication by ${\mathbb{Q}}(\sqrt{-15})$, which is not contained in $H$. Furthermore, the number field LMFDB data shows that $\ker{\tilde{\rho}_{g}}\subset G_{H}$, so $\tilde{\rho}_{g}(G_{{\mathbb{Q}}}):\tilde{\rho}_{g}(G_{H})]=[\rho_{g}(G_{{\mathbb{Q}}}):\rho_{g}(G_{H})]=2$. Since $a_{139}(g)=\pm 2i$, $iI_{2}\in\rho_{g}(G_{H})\subset\rho_{g}(G_{{\mathbb{Q}}})$. Since the coefficient field of $g$ is ${\mathbb{Q}}(i,\sqrt{6})$, it does not contain a primitive $8$-th root of unity, so the scalar matrices of $\rho_{g}(G_{{\mathbb{Q}}})$ are exactly $\pm iI_{2},\pm I_{2}$. Finally, $\varepsilon_{g}(G_{H})=\varepsilon_{g}(G_{{\mathbb{Q}}})=\{\pm i,\pm 1\}$, since $\varepsilon_{g}$ and $H$ have coprime conductors.